# The metaplectic semigroup and its applications to time-frequency analysis and evolution operators

### v1

@@ 4.3. Intertwining with the Schrödinger representation
For the rest of this section, we denote by $\varphi_{M}(t)=e^{-\pi Mt\cdot t}$, where $M\in\mathbb{R}^{d\times d}$ is symmetric and positive-definite.
@@ 4.3.1. Complex translations and modulations
(151) $M_{w}f(t)=e^{2\pi iw\cdot t}f(t),\qquad t\in\mathbb{R}^{d},\quad w\in\mathbb{C}^{d}.$
@@
(152) $T_{z}f(t)=\int_{\mathbb{R}^{d}}\widehat{f}(\eta)e^{2\pi\eta\cdot(t-z)}d\eta,\qquad t\in\mathbb{R}^{d}.$
@@
(153) $\rho(z,w;\tau)f(t)=e^{2\pi i\tau}e^{-i\pi z\cdot w}e^{2\pi iw\cdot t}\int_{\mathbb{R}^{d}}\widehat{f}(\eta)e^{2\pi i\eta\cdot(t-z)}d\eta,\qquad t\in\mathbb{R}^{d}.$
@@
(154) $\int_{\mathbb{R}^{d}}\widehat{\varphi_{M}}(\eta)e^{2\pi i\eta\cdot(t-z)}d\eta=e^{-\pi M(t-z)\cdot(t-z)},$
@@
(155) $\rho(z,w;\tau)\varphi_{M}(t)=e^{2\pi i\tau}e^{-i\pi z\cdot w}e^{2\pi iw\cdot t}e^{-\pi M(t-z)(t-z)},\qquad t\in\mathbb{R}^{d}.$
@@
Clearly, it is enough to prove the assertion for the generators of $\operatorname*{Mp}(d,\mathbb{R})$. First, we consider the Fourier transform:
(157) $\mathcal{F}\rho(z,w;\tau)\varphi_{M}(t)$ $=i^{-d/2}\int_{\mathbb{R}^{d}}\rho(z,w;\tau)\varphi_{M}(y)e^{-2\pi iy\cdot t}dy$
(158) $=i^{-d/2}e^{2\pi i\tau}e^{-i\pi z\cdot w}\int_{\mathbb{R}^{d}}e^{2\pi iw\cdot y}e^{-\pi M(y-z)\cdot(y-z)}e^{-2\pi it\cdot y}dy$
(159) $=i^{-d/2}e^{2\pi i\tau}e^{-i\pi z\cdot w}\int_{\mathbb{R}^{d}}e^{2\pi iw\cdot(u+z)}e^{-\pi Mu\cdot u}e^{-2\pi it\cdot(u+z)}du$
(160) $=i^{-d/2}e^{2\pi i\tau}e^{i\pi z\cdot w}e^{-2\pi it\cdot z}\int_{\mathbb{R}^{d}}e^{-\pi Mu\cdot u}e^{-2\pi iu\cdot(t-w)}du$
(161) $=\rho(w,-z;\tau)\mathcal{F}\varphi_{M}(t)=\rho(J(z,w);\tau)\mathcal{F}\varphi_{M}(t).$
@@
(162) $\mathfrak{T}_{E}\rho(z,w;\tau)\varphi_{M}(t)$ $=i^{m}|\det(E)|^{1/2}[\rho(z,w;\tau)\varphi_{M}](Et)$
(163) $=i^{m}|\det(E)|^{1/2}e^{2\pi i\tau}e^{-i\pi z\cdot w}e^{2\pi iEt\cdot w}e^{-\pi M(Et-z)\cdot(Et-z)}$
(164) $=i^{m}|\det(E)|^{1/2}e^{2\pi i\tau}e^{-i\pi(EE^{-1})z\cdot w}e^{2\pi it\cdot E^{\top}w}e^{-\pi M[E(t-E^{-1}z)]\cdot[E(t-E^{-1}z)]}$
(165) $=i^{m}|\det(E)|^{1/2}e^{2\pi i\tau}e^{-i\pi E^{-1}z\cdot E^{\top}w}e^{2\pi it\cdot E^{\top}w}e^{-\pi M[E(t-E^{-1}z)]\cdot[E(t-E^{-1}z)]}$
(166) $=i^{m}|\det(E)|^{1/2}e^{2\pi i\tau}e^{-i\pi E^{-1}z\cdot E^{\top}w}e^{2\pi it\cdot E^{\top}w}e^{-\pi M[E(t-E^{-1}z)]\cdot[E(t-E^{-1}z)]}$
(167) $=\rho(E^{-1}z,E^{\top}w;\tau)\mathfrak{T}_{E}\varphi_{M}(t)=\rho(\mathcal{D}_{E}(z,w);\tau)\mathfrak{T}_{E}\varphi_{M}(t),$
@@
(168) $\begin{split}\rho(V_{Q}(z,w);\tau)\mathfrak{p}_{Q}\varphi_{M}(t)&=\rho(z,w+Qz;\tau)\mathfrak{p}_{Q}\varphi_{M}(t)\\ &=e^{2\pi i\tau}e^{-i\pi z\cdot(w+Qz)}e^{2\pi i(w+Qz)\cdot t}e^{i\pi Q(t-z)\cdot(t-z)}e^{-\pi M(t-z)\cdot(t-z)}\\ &=e^{i\pi Qt\cdot t}e^{2\pi i\tau}e^{-i\pi z\cdot w}e^{2\pi iw\cdot t}e^{-\pi M(t-z)\cdot(t-z)}\\ &=\mathfrak{p}_{Q}\rho(z,w;\tau)\varphi_{M}(t).\end{split}$
@@
when $\varphi_{a}(t)=e^{-\pi at^{2}}$, $(z,w)=(x,\xi)\in\mathbb{R}^{2}$ and $Q=i$. Let $\vartheta>0$, we show that
@@
Therefore $\mathfrak{R}_{\vartheta}\varphi_{a}$ is a Gaussian and $T_{z}\mathfrak{R}_{\vartheta}\varphi_{a}(t)=\mathfrak{R}_{\vartheta}\varphi_{a}(t-z)$ for every $z\in\mathbb{C}$. Then, by using the identity $\cosh\vartheta^{2}-\sinh\vartheta^{2}=1$ and its surrogates, we obtain
(173) $\mathfrak{R}_{\vartheta}$ $\rho(x,\xi;\tau)\varphi_{a}(t)=(\cosh\vartheta)^{-1/2}\int_{-\infty}^{\infty}(\rho(x,\xi;\tau)\varphi_{a}){\text{\textasciicircum}}(\eta)e^{i\pi(2t\eta+i(t^{2}+\eta^{2})\sinh\vartheta)/\cosh\vartheta}d\eta$
(174) $=(\cosh\vartheta)^{-1/2}\int_{-\infty}^{\infty}\rho(\xi,-x;\tau)\widehat{\varphi_{a}}(\eta)e^{i\pi(2t\eta+i(t^{2}+\eta^{2})\sinh\vartheta)/\cosh\vartheta}d\eta$
(175) $=(\cosh\vartheta)^{-1/2}e^{2\pi i\tau}e^{i\pi x\cdot\xi}\int_{-\infty}^{\infty}e^{-2\pi i\eta\cdot x}\widehat{\varphi_{a}}(\eta-\xi)e^{i\pi(2t\eta+i(t^{2}+\eta^{2})\sinh\vartheta)/\cosh\vartheta}d\eta$
(176) $=(\cosh\vartheta)^{-1/2}e^{2\pi i\tau}e^{i\pi x\cdot\xi}\int_{-\infty}^{\infty}e^{-2\pi i(\omega+\xi)\cdot x}\widehat{\varphi_{a}}(\omega)e^{i\pi(2t(\omega+\xi)+i(t^{2}+(\omega+\xi)^{2})\sinh\vartheta)/\cosh\vartheta}d\eta$
(177) $=(\cosh\vartheta)^{-1/2}e^{2\pi i\tau}e^{-i\pi x\cdot\xi}e^{-\pi\tanh\vartheta t^{2}}e^{2\pi it\xi/\cosh\vartheta}$
(178) $\qquad\qquad\qquad\qquad\qquad\qquad\qquad\times\int_{-\infty}^{\infty}e^{-2\pi i\omega\cdot x}\widehat{\varphi_{a}}(\omega)e^{i\pi(2t\omega+i(\omega+\xi)^{2}\sinh\vartheta)/\cosh\vartheta}d\omega$
(179) $=(\cosh\vartheta)^{-1/2}e^{2\pi i\tau}e^{-i\pi x\cdot\xi}e^{-\pi\tanh\vartheta t^{2}}e^{2\pi it\xi/\cosh\vartheta}e^{-\pi\tanh\vartheta\xi^{2}}\times$
(180) $\qquad\qquad\qquad\qquad\qquad\qquad\qquad\times\int_{-\infty}^{\infty}\widehat{\varphi_{a}}(\omega)e^{2\pi i\omega(t-x\cosh\vartheta+i\xi\sinh\vartheta)/\cosh\vartheta}e^{-\pi\tanh\vartheta\omega^{2}}d\omega$
(181) $=(\cosh\vartheta)^{-1/2}e^{2\pi i\tau}e^{-i\pi x\cdot\xi}e^{-\pi\tanh\vartheta t^{2}}e^{-\pi\tanh\vartheta\xi^{2}}e^{2\pi it\xi/\cosh\vartheta}e^{\pi\tanh\vartheta(t-x\cosh\vartheta+i\xi\sinh\vartheta)^{2}}\times$
(182) $\times\underbrace{\int_{-\infty}^{\infty}\widehat{\varphi_{a}}(\omega)e^{2\pi i\omega(t-x\cosh\vartheta+i\xi\sinh\vartheta)/\cosh\vartheta}e^{-\pi\tanh\vartheta(\omega^{2}+(t-x\cosh\vartheta+i\xi\sinh\vartheta)^{2})}d\omega}_{=\mathfrak{R}_{\vartheta}\varphi_{a}(t-x\cosh\vartheta+i\xi\sinh\vartheta)=T_{x\cosh\vartheta-i\xi\sinh\vartheta}\mathfrak{R}_{\vartheta}\varphi_{a}(t)}$
(183) $=e^{2\pi i\tau}e^{-i\pi(x\cosh\vartheta-i\xi\sinh\vartheta)(ix\sinh\vartheta+\xi\cosh\vartheta)}e^{2\pi it(ix\sinh\vartheta+\xi\cosh\vartheta)}T_{x\cosh\vartheta-i\xi\sinh\vartheta}\mathfrak{R}_{\vartheta}\varphi_{a}(t)$
(184) $=\rho(x\cosh\vartheta-i\xi\sinh\vartheta,ix\sinh\vartheta+\xi\cosh\vartheta;\tau)\mathfrak{R}_{\vartheta}\varphi_{a}(t).$
@@
Claim 2. The assertion holds in dimension $d\geq 1$ for Gaussian states $\varphi_{\Delta}(t)=e^{-\pi\Delta t\cdot t}$, $\Delta>0$ diagonal.
@@
for Gaussians $f_{1},\ldots,f_{d}$, $x,\xi\in\mathbb{R}^{d}$ and $\tau\in\mathbb{R}$, and $S_{1},\ldots,S_{d}\in\operatorname*{Sp}_{+}(d,\mathbb{C})$. In particular, if $\Xi$ is purely positive, $\Xi=\Xi_{1}\otimes\ldots\otimes\Xi_{d}$, with either $\Xi_{j}=V_{i}$ or $\Xi_{j}=\mathcal{R}_{\vartheta_{j}}$ ($\vartheta_{j}\geq 0$), $j=1,\ldots,d$, we have
@@
(188) $\widehat{\Xi}\rho(x,\xi;\tau)\varphi_{\Delta}(t)$ $=e^{-2\pi i\tau(d-1)}\Big(\bigotimes_{j=1}^{d}\widehat{\Xi}_{j}\bigotimes_{j=1}^{d}\rho(x_{j},\xi_{j};\tau)\Big)\varphi_{\Delta}(t)$
(189) $=e^{-2\pi i\tau(d-1)}\bigotimes_{j=1}^{d}\widehat{\Xi}_{j}\rho(x_{j},\xi_{j};\tau)\varphi_{a_{j}}(t_{j})$
(190) $=e^{-2\pi i\tau(d-1)}\bigotimes_{j=1}^{d}\rho(\Xi_{j}(x_{j},\xi_{j});\tau)\widehat{\Xi}_{j}\varphi_{a_{j}}(t_{j})$
(191) $=e^{-2\pi i\tau(d-1)}\rho\Big(\bigotimes_{j=1}^{d}\Xi_{j}(x_{j},\xi_{j});\tau\Big)\bigotimes_{j=1}^{d}\widehat{\Xi}_{j}\varphi_{a_{j}}(t_{j})$
(192) $=\rho(\Xi(x,\xi);\tau)\widehat{\Xi}\varphi_{\Delta}(t).$
@@
Finally, Theorem 4.6 follows by combining Lemmas 4.8 and 4.9. To be precise, let $\widehat{S}\in\operatorname*{Mp}_{+}(d,\mathbb{C})$ and observe that both sides of (150) are well-defined in $L^{2}(\mathbb{R}^{d})$. By item $(ii)$ of Theorem 3.10, we can write $\widehat{S}=\widehat{U}_{1}\widehat{\Xi}\widehat{U}_{2}$, where $\widehat{U}_{1},\widehat{U}_{2}\in\operatorname*{Mp}(d,\mathbb{R})$ and $\widehat{\Xi}$ is an atomic metaplectic contraction. Then, by applying (54), Lemma 4.9 and Lemma 4.8 in the order,
@@
Consequently, (150) holds for functions in the form (198) with $N=1$, and the assertion follows by linearity. This proves $(i)$.

### v2

@@ 4.3.1. Complex translations and modulations
(151) $M_{w}f(y)=e^{2\pi iw\cdot y}f(y),\qquad y\in\mathbb{R}^{d},\quad w\in\mathbb{C}^{d}.$
@@
(152) $T_{z}f(y)=\int_{\mathbb{R}^{d}}\widehat{f}(\eta)e^{2\pi i\eta\cdot(y-z)}d\eta,\qquad y\in\mathbb{R}^{d}.$
@@
(153) $\rho(z,w;\tau)f(y)=e^{2\pi i\tau}e^{-i\pi z\cdot w}e^{2\pi iw\cdot y}\int_{\mathbb{R}^{d}}\widehat{f}(\eta)e^{2\pi i\eta\cdot(y-z)}d\eta,\qquad y\in\mathbb{R}^{d}.$
@@
(154) $\int_{\mathbb{R}^{d}}\widehat{\varphi_{M}}(\eta)e^{2\pi i\eta\cdot(y-z)}d\eta=e^{-\pi M(y-z)\cdot(y-z)},$
@@
(155) $\rho(z,w;\tau)\varphi_{M}(y)=e^{2\pi i\tau}e^{-i\pi z\cdot w}e^{2\pi iw\cdot y}e^{-\pi M(y-z)(y-z)},\qquad y\in\mathbb{R}^{d}.$
@@
Clearly, it is enough to prove the assertion for the generators of $\operatorname*{Mp}(d,\mathbb{R})$. First, we consider the Fourier transform. By Proposition 2.2,
(157) $\mathcal{F}\rho(z,w;\tau)\varphi_{M}(x)$ $=i^{-d/2}\int_{\mathbb{R}^{d}}\rho(z,w;\tau)\varphi_{M}(y)e^{-2\pi iy\cdot x}dy$
(158) $=i^{-d/2}e^{2\pi i\tau}e^{-i\pi z\cdot w}\int_{\mathbb{R}^{d}}e^{2\pi iw\cdot y}e^{-\pi M(y-z)\cdot(y-z)}e^{-2\pi ix\cdot y}dy$
(159) $=i^{-d/2}e^{2\pi i\tau}e^{-i\pi z\cdot w}e^{-\pi Mz\cdot z}\int_{\mathbb{R}^{d}}e^{-\pi My\cdot y}e^{-2\pi iy\cdot(x-w+iMz)}dy$
(160) $=i^{-d/2}\det(M)^{-1/2}e^{2\pi i\tau}e^{-i\pi z\cdot w}e^{-\pi Mz\cdot z}e^{-\pi M^{-1}(x-w+iMz)\cdot(x-w+iMz)}$
(161) $=i^{-d/2}\det(M)^{-1/2}e^{2\pi i\tau}e^{-i\pi z\cdot w}e^{-\pi M^{-1}(x-w)\cdot(x-w)}e^{-2\pi i(x-w)\cdot z}$
(162) $=i^{-d/2}e^{2\pi i\tau}e^{i\pi z\cdot w}e^{-2\pi ix\cdot z}\mathcal{F}\varphi_{M}(x-w)$
(163) $=\rho(w,-z;\tau)\mathcal{F}\varphi_{M}(x)=\rho(J(z,w);\tau)\mathcal{F}\varphi_{M}(x).$
@@
(164) $\mathfrak{T}_{E}\rho(z,w;\tau)\varphi_{M}(x)$ $=i^{m}|\det(E)|^{1/2}[\rho(z,w;\tau)\varphi_{M}](Ex)$
(165) $=i^{m}|\det(E)|^{1/2}e^{2\pi i\tau}e^{-i\pi z\cdot w}e^{2\pi iEx\cdot w}e^{-\pi M(Ex-z)\cdot(Ex-z)}$
(166) $=i^{m}|\det(E)|^{1/2}e^{2\pi i\tau}e^{-i\pi(EE^{-1})z\cdot w}e^{2\pi ix\cdot E^{\top}w}e^{-\pi M[E(x-E^{-1}z)]\cdot[E(x-E^{-1}z)]}$
(167) $=i^{m}|\det(E)|^{1/2}e^{2\pi i\tau}e^{-i\pi E^{-1}z\cdot E^{\top}w}e^{2\pi ix\cdot E^{\top}w}e^{-\pi M[E(x-E^{-1}z)]\cdot[E(x-E^{-1}z)]}$
(168) $=i^{m}|\det(E)|^{1/2}e^{2\pi i\tau}e^{-i\pi E^{-1}z\cdot E^{\top}w}e^{2\pi ix\cdot E^{\top}w}e^{-\pi M[E(x-E^{-1}z)]\cdot[E(x-E^{-1}z)]}$
(169) $=\rho(E^{-1}z,E^{\top}w;\tau)\mathfrak{T}_{E}\varphi_{M}(x)=\rho(\mathcal{D}_{E}(z,w);\tau)\mathfrak{T}_{E}\varphi_{M}(x),$
@@
(170) $\begin{split}\rho(V_{Q}(z,w);\tau)\mathfrak{p}_{Q}\varphi_{M}(x)&=\rho(z,w+Qz;\tau)\mathfrak{p}_{Q}\varphi_{M}(x)\\ &=e^{2\pi i\tau}e^{-i\pi z\cdot(w+Qz)}e^{2\pi i(w+Qz)\cdot x}e^{i\pi Q(x-z)\cdot(x-z)}e^{-\pi M(x-z)\cdot(x-z)}\\ &=e^{i\pi Qx\cdot x}e^{2\pi i\tau}e^{-i\pi z\cdot w}e^{2\pi iw\cdot x}e^{-\pi M(x-z)\cdot(x-z)}\\ &=\mathfrak{p}_{Q}\rho(z,w;\tau)\varphi_{M}(x).\end{split}$
@@
when $\varphi_{a}(y)=e^{-\pi ay^{2}}$, $(z,w)=(x,\xi)\in\mathbb{R}^{2}$ and $Q=i$. Let $\vartheta>0$, we show that
@@
Therefore $\mathfrak{R}_{\vartheta}\varphi_{a}$ is a Gaussian and $T_{z}\mathfrak{R}_{\vartheta}\varphi_{a}(y)=\mathfrak{R}_{\vartheta}\varphi_{a}(y-z)$ for every $z\in\mathbb{C}$. Then, by using the identity $\cosh\vartheta^{2}-\sinh\vartheta^{2}=1$ and its surrogates, we obtain
(175) $\mathfrak{R}_{\vartheta}$ $\rho(x,\xi;\tau)\varphi_{a}(y)=(\cosh\vartheta)^{-1/2}\int_{-\infty}^{\infty}(\rho(x,\xi;\tau)\varphi_{a}){\text{\textasciicircum}}(\eta)e^{i\pi(2y\eta+i(y^{2}+\eta^{2})\sinh\vartheta)/\cosh\vartheta}d\eta$
(176) $=(\cosh\vartheta)^{-1/2}\int_{-\infty}^{\infty}\rho(\xi,-x;\tau)\widehat{\varphi_{a}}(\eta)e^{i\pi(2y\eta+i(y^{2}+\eta^{2})\sinh\vartheta)/\cosh\vartheta}d\eta$
(177) $=(\cosh\vartheta)^{-1/2}e^{2\pi i\tau}e^{i\pi x\cdot\xi}\int_{-\infty}^{\infty}e^{-2\pi i\eta\cdot x}\widehat{\varphi_{a}}(\eta-\xi)e^{i\pi(2y\eta+i(y^{2}+\eta^{2})\sinh\vartheta)/\cosh\vartheta}d\eta$
(178) $=(\cosh\vartheta)^{-1/2}e^{2\pi i\tau}e^{i\pi x\cdot\xi}\int_{-\infty}^{\infty}e^{-2\pi i(\omega+\xi)\cdot x}\widehat{\varphi_{a}}(\omega)e^{i\pi(2y(\omega+\xi)+i(y^{2}+(\omega+\xi)^{2})\sinh\vartheta)/\cosh\vartheta}d\eta$
(179) $=(\cosh\vartheta)^{-1/2}e^{2\pi i\tau}e^{-i\pi x\cdot\xi}e^{-\pi\tanh\vartheta y^{2}}e^{2\pi iy\xi/\cosh\vartheta}$
(180) $\qquad\qquad\qquad\qquad\qquad\qquad\qquad\times\int_{-\infty}^{\infty}e^{-2\pi i\omega\cdot x}\widehat{\varphi_{a}}(\omega)e^{i\pi(2y\omega+i(\omega+\xi)^{2}\sinh\vartheta)/\cosh\vartheta}d\omega$
(181) $=(\cosh\vartheta)^{-1/2}e^{2\pi i\tau}e^{-i\pi x\cdot\xi}e^{-\pi\tanh\vartheta y^{2}}e^{2\pi iy\xi/\cosh\vartheta}e^{-\pi\tanh\vartheta\xi^{2}}\times$
(182) $\qquad\qquad\qquad\qquad\qquad\qquad\qquad\times\int_{-\infty}^{\infty}\widehat{\varphi_{a}}(\omega)e^{2\pi i\omega(y-x\cosh\vartheta+i\xi\sinh\vartheta)/\cosh\vartheta}e^{-\pi\tanh\vartheta\omega^{2}}d\omega$
(183) $=(\cosh\vartheta)^{-1/2}e^{2\pi i\tau}e^{-i\pi x\cdot\xi}e^{-\pi\tanh\vartheta y^{2}}e^{-\pi\tanh\vartheta\xi^{2}}e^{2\pi iy\xi/\cosh\vartheta}e^{\pi\tanh\vartheta(y-x\cosh\vartheta+i\xi\sinh\vartheta)^{2}}\times$
(184) $\times\underbrace{\int_{-\infty}^{\infty}\widehat{\varphi_{a}}(\omega)e^{2\pi i\omega(y-x\cosh\vartheta+i\xi\sinh\vartheta)/\cosh\vartheta}e^{-\pi\tanh\vartheta(\omega^{2}+(y-x\cosh\vartheta+i\xi\sinh\vartheta)^{2})}d\omega}_{=\mathfrak{R}_{\vartheta}\varphi_{a}(y-x\cosh\vartheta+i\xi\sinh\vartheta)=T_{x\cosh\vartheta-i\xi\sinh\vartheta}\mathfrak{R}_{\vartheta}\varphi_{a}}$
(185) $=e^{2\pi i\tau}e^{-i\pi(x\cosh\vartheta-i\xi\sinh\vartheta)(ix\sinh\vartheta+\xi\cosh\vartheta)}e^{2\pi iy(ix\sinh\vartheta+\xi\cosh\vartheta)}T_{x\cosh\vartheta-i\xi\sinh\vartheta}\mathfrak{R}_{\vartheta}\varphi_{a}(y)$
(186) $=\rho(x\cosh\vartheta-i\xi\sinh\vartheta,ix\sinh\vartheta+\xi\cosh\vartheta;\tau)\mathfrak{R}_{\vartheta}\varphi_{a}(y).$
@@
Claim 2. The assertion holds in dimension $d\geq 1$ for Gaussian states $\varphi_{\Delta}(y)=e^{-\pi\Delta y\cdot y}$, $\Delta>0$ diagonal.
@@
for Gaussians $f_{1},\ldots,f_{d}$, $x,\xi\in\mathbb{R}^{d}$ and $\tau\in\mathbb{R}$, and $S_{1},\ldots,S_{d}\in\operatorname*{Sp}_{+}(1,\mathbb{C})$. In particular, if $\Xi$ is purely positive, $\Xi=\Xi_{1}\otimes\ldots\otimes\Xi_{d}$, with either $\Xi_{j}=V_{i}$ or $\Xi_{j}=\mathcal{R}_{\vartheta_{j}}$ ($\vartheta_{j}\geq 0$), $j=1,\ldots,d$, we have
@@
(190) $\widehat{\Xi}\rho(x,\xi;\tau)\varphi_{\Delta}(y)$ $=e^{-2\pi i\tau(d-1)}\Big(\bigotimes_{j=1}^{d}\widehat{\Xi}_{j}\bigotimes_{j=1}^{d}\rho(x_{j},\xi_{j};\tau)\Big)\varphi_{\Delta}(y)$
(191) $=e^{-2\pi i\tau(d-1)}\bigotimes_{j=1}^{d}\widehat{\Xi}_{j}\rho(x_{j},\xi_{j};\tau)\varphi_{a_{j}}(y_{j})$
(192) $=e^{-2\pi i\tau(d-1)}\bigotimes_{j=1}^{d}\rho(\Xi_{j}(x_{j},\xi_{j});\tau)\widehat{\Xi}_{j}\varphi_{a_{j}}(y_{j})$
(193) $=e^{-2\pi i\tau(d-1)}\rho\Big(\bigotimes_{j=1}^{d}\Xi_{j}(x_{j},\xi_{j});\tau\Big)\bigotimes_{j=1}^{d}\widehat{\Xi}_{j}\varphi_{a_{j}}(y_{j})$
(194) $=\rho(\Xi(x,\xi);\tau)\widehat{\Xi}\varphi_{\Delta}(y).$
@@
In conclusion the proof of Theorem 4.6 follows by combining Lemmas 4.8 and 4.9. To be precise, let $\widehat{S}\in\operatorname*{Mp}_{+}(d,\mathbb{C})$ and observe that both sides of (150) are well-defined in $L^{2}(\mathbb{R}^{d})$. By item $(ii)$ of Theorem 3.10, we can write $\widehat{S}=\widehat{U}_{1}\widehat{\Xi}\widehat{U}_{2}$, where $\widehat{U}_{1},\widehat{U}_{2}\in\operatorname*{Mp}(d,\mathbb{R})$ and $\widehat{\Xi}$ is an atomic metaplectic contraction. Then, by applying (54), Lemma 4.9 and Lemma 4.8 in the order,
@@
Consequently, (150) holds for functions in the form (200) with $N=1$, and the assertion follows by linearity.